40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 26 17:05:44 2009, rob, Update, PSL, MOPA re-activated DSC_0513.JPGDSC_0517.JPG
    Reply  Tue May 26 17:34:14 2009, rob, Update, PSL, MOPA re-deactivated laser_temp.jpg
       Reply  Wed May 27 10:54:09 2009, rob, Update, PSL, we don't understand the chiller (broken) 
          Reply  Wed May 27 15:59:44 2009, rob, Update, PSL, we don't understand the chiller (broken) 
             Reply  Thu May 28 14:34:25 2009, rob, Update, PSL, chiller diagnosis 
Message ID: 1629     Entry time: Thu May 28 14:34:25 2009     In reply to: 1628
Author: rob 
Type: Update 
Category: PSL 
Subject: chiller diagnosis 

Quote:

steve, alberto, rob

After some futzing around with the chiller, we have come to the tentative conclusion that the refrigeration unit is not working.  Steve called facilities to try to get them to recharge the refrigerant (R-404a) tomorrow, and we're also calling around for a spare chiller somewhere in the project (without luck so far).

 The repair man thinks it's a bad start capacitor, which is 240uF at 120V.  Steve has ordered a new one which should be here tomorrow, and with luck we'll have lasing by tomorrow afternoon.

ELOG V3.1.3-